OpenStudy (blake57roger):

Find the exact area of the circle with the given radius. r = 4.5 m

OpenStudy (mathstudent55):

Do you know the formula for the area of a circle?

OpenStudy (mathstudent55):

The area of a circle is calculated according to the formula below: \(A = \pi r^2\) where A = area \(\pi\) is a constant number that is approximately 3.14 r = radius of the circle

OpenStudy (mathstudent55):

Your problem wants the exact area, so use the symbol \(\pi\) in your answer. All you need to do is use the formula, and replace r with its value, 4.5 m.

OpenStudy (mathstudent55):

\(A = \pi \times (4.5 ~m)^2\)